Help! All four are incorrect!

A 15 kg block is accelerated at 1.6 m/s2 up a frictionless plane inclined at 5.0
Answer #1

time taken to reach the speed 36 m / s from 10 m/ s is t = ( v - u ) / a
                                                                                     = 16.25 s
(a) . the change in the block's mechanical energy = ( 1/ 2) m[v^ 2- u^ 2]
                                                                          = 8970 J
                                                                      E =8970 kJ
(b) the average rate at which energy is transferred to the block P = E / t
                   P = 552 W
(c). instantaneous rate of that transfer when the block's speed is 10 m/s is P ' = Fu
where F = ma
So, P ' = mau = W
(d)  instantaneous rate of that transfer when the block's speed is 36 m/s is P " = Fv
        P" = mav
            = W